OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form.

OpenStudy (anonymous):

point slope form: (y-y1)=Slope(x-x1) sub y1=4 and x=-3 and slope=2 y-4=2(x-(-3)) simplify it y-4=2x+6 y=2x+10

OpenStudy (anonymous):

point slope form is y=mx+c m=2 so y=2x+c it passes through (-3,4) put x=-3 and y=4 4=2*-3+c 4=-6+c c=10 so y=2x+10:)

OpenStudy (anonymous):

Thank you! do you guys know "Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form"

OpenStudy (anonymous):

 :) medal please:)lol